# Understanding and Evaluating Human Preferences for AI Generated Images with Instruction Tuning

### v1

@@ Abstract
Artificial Intelligence Generated Content (AIGC) has grown rapidly in recent years, among which AI-based image generation has gained widespread attention due to its efficient and imaginative image creation ability. However, AI-generated Images (AIGIs) may not satisfy human preferences due to their unique distortions, which highlights the necessity to understand and evaluate human preferences for AIGIs. To this end, in this paper, we first establish a novel Image Quality Assessment (IQA) database for AIGIs, termed AIGCIQA2023+, which provides human visual preference scores and detailed preference explanations from three perspectives including quality, authenticity, and correspondence. Then, based on the constructed AIGCIQA2023+ database, this paper presents a MINT-IQA model to evaluate and explain human preferences for AIGIs from Multi-perspectives with INstruction Tuning. Specifically, the MINT-IQA model first learn and evaluate human preferences for AI-generated Images from multi-perspectives, then via the vision-language instruction tuning strategy, MINT-IQA attains powerful understanding and explanation ability for human visual preference on AIGIs, which can be used for feedback to further improve the assessment capabilities. Extensive experimental results demonstrate that the proposed MINT-IQA model achieves state-of-the-art performance in understanding and evaluating human visual preferences for AIGIs, and the proposed model also achieves competing results on traditional IQA tasks compared with state-of-the-art IQA models. The AIGCIQA2023+ database and MINT-IQA model will be released to facilitate future research on https://github.com/IntMeGroup/MINT-IQA
@@ I Introduction
Artificial Intelligence Generated Content (AIGC) refers to the content generated with the assistance of AI, including texts, images, audios, etc. As an important part of AIGC, AI Generated Images (AIGIs) have gained significant attention due to their broad application prospects, thus, many AI-based image generation methods have been proposed in recent years, such as DALLE [1], Stable-diffusion [2], Unidiffuser [3], etc. However, due to the limitations of model capacity, computing overhead, and human-AI interaction gap, etc., not all generated images can satisfy users’ requirements, and many AIGIs do not conform to human visual preferences or expectations. Unlike common image content, such as Natural Scene Images (NSIs) [4], Screen Content Images (SCIs) [5] etc., which generally encounters common degradations such as noise, blur, compression [6, 7, 8], AIGIs exhibit unique distortions including unreal structures, unreasonable component combinations, etc. [9, 10, 11] Moreover, for text-to-image generation methods, the generated images may not accurately correspond to the semantics of the text prompts or may not meet the expectations of users [12, 13].
@@
To address the aforementioned challenges, in this paper, we first extend the previous AIGCIQA2023 database [17] to AIGCIQA2023+, towards better understanding and evaluating human preferences for AIGIs, and then propose a novel method, termed MINT-IQA, towards automatically evaluating and explaining the visual preference and expectation attributes for AIGIs from Multi-perspectives with INstruction Tuning. Specifically, as shown in Fig. 1, the established AIGCIQA2023+ contains 2400 AIGIs generated using 6 AIGI models based on 100 text prompts, 7,200 Mean Opinion Scores (MOSs) and 16,800 detailed preference-explanation annotations for these AIGIs. The evaluation perspectives include quality, authenticity, and text-image correspondence. We further propose a MINT-IQA model, which allows for a more comprehensive evaluation of the human preferences for AIGIs from multi-perspectives, resulting in state-of-the-art performance on AIGC IQA databases. To further enhance the degradation understanding and explanation capabilities of our model, we adopt the vision-language instruction tuning strategy, and fine-tune the MINT-IQA model based on the AIGCIQA2023+ database. This approach effectively leverages the frozen Large Language Model (LLM) to provide detailed explanations of the human preference basis for each AIGI as shown in Fig. 2, which can be used for feedback to further improve the assessment capabilities as shown in Section V. Additionally, MINT-IQA model also achieves the best performance on traditional IQA databases, which further demonstrates its generality in understanding and evaluating human preferences.
@@
  • We establish an AI generated image quality assessment database, termed AIGCIQA2023+, to conduct comprehensive understanding and evaluation research on human visual preferences for AIGIs.
@@ II-A Image Quality Assessment Databases
Traditional IQA Databases CID2013 [22] Rating Overall 480 14,880 No
@@
Many IQA databases have been established in the literature [27, 28, 29]. As shown in Table I, according to the image acquisition method, they can be categorized into two types: (1) traditional IQA databases and (2) AIGC IQA databases. Traditional IQA databases consist of natural images with realistic camera distortions such as noise, blur, compression, etc. CID2013 [22] is a small database consists of 585 realistically blurred pictures taken by 79 cameras. KonIQ-10k [30] and SPAQ [24] expand the data scale and cover a wide range of scene categories. AVA [25] is a large-scale database constructed for aesthetic image quality assessment. Recently, the development of text-to-image generative models have led to an explosion of AI-generated images. Thus many AIGC IQA databases have also been constructed, which contain generated images with unique distortions including unreal structures, unreasonable component combinations, etc. DiffusionDB [26] is a large-scale database containing over 1.8 million text-image pairs without score-specific annotations. Pick-A-Pic [14] and HPS [9] provide subjective annotations of human preferences by selecting the most preferred one among a group or a pair of AIGIs. ImageReward [15] contains preference ranking annotations for AIGIs. The aforementioned AIGC IQA databases only evaluate the AIGIs from the overall dimension, which cannot reflect and characterize complex human visual preferences for AIGIs. AGIQA-3K [16] refines the rating granularity by scoring two dimensions including perception and alignment. Our preliminary AIGCIQA2023 database [17] contains subjective quality ratings from three dimensions including quality, authenticity, and correspondence. However, the aforementioned works lack detailed annotations of the human preference for AIGIs, which motivates us to further extend the AIGCIQA2023 database to include language annotations.
@@ II-B Image Quality Assessment Models
Image quality assessment algorithms aim to objectively and quantitatively evaluate the perceptual quality of images. Based on different feature extraction and evaluation methods, current popular models for image quality assessment can be classified into three categories: (1) Handcrafted-based models, including: NIQE [31], BRISQUE [32], QAC [33], BMPRI [34], HOSA [35], BPRI [36], HIGRADE [37], etc. These models generally extract features based on prior knowledge related to image quality. (2) Deep learning-based models, including: CNNIQA [38], DBCNN [39], WaDIQaM-NR [40], StairIQA [41], etc. These models characterize quality-aware information by training deep neural networks using labeled data. (3) With the popularity of text-to-image generation, many vision-language pre-training models have been adopted to evaluate the text-image alignment, including: CLIP [20], BLIP [21], FLIP [42], Aesthetic [43], HPS [9], PickScore [14], ImageReward [15], StairReward [16], etc. These models are often used for evaluating generated images based on its text prompt considering a mixture of elements such as text-image alignment, fidelity, aesthetics, etc.
The aforementioned models can only evaluate a single-dimension preference for AIGIs, which lacks the ability to conduct a comprehensive assessment from multiple perspectives. This motivates us to propose a unified model to evaluate AIGIs from multiple dimensions.
@@ III-A AIGI Collection
We adopt six latest text-to-image generative models, including Glide[44], Lafite[45], DALLE[46], Stable-diffusion[47], Unidiffuser[3], Controlnet[48], to produce AIGIs by using open source code and default weights. To ensure content diversity and catch up with the practical application requirements, we collect diverse texts from the PartiPrompts website [49] as the prompts for AI-based image generation. The text prompts can be simple, allowing generative models to produce imaginative results. They can also be complex, which raises the challenge for generative models. We select 10 scene categories from the prompt set, and each scene contains 10 challenge categories. Overall, we collect 100 text prompts (10 scene categories $\times$ 10 challenge categories) from PartiPrompts[49]. Based on the selected prompts, we generate 4 different images for each generative model. Therefore, the constructed database totally contains 2400 AIGIs (4 images $\times$ 6 models $\times$ 100 prompts).
@@ III-B Subjective Experiment Setup
The AIGIs are presented in a random order on an iMac monitor with a resolution of 4096 × 2304, using an interface designed with Python Tkinter as shown in Fig. 4. The interface allows viewers to browse the previous and next AIGIs, and evaluate them by scoring within the range from 0 to 5 as well as providing fine-grained explanations. A total of 76 graduate students participate in the experiment. The subjects are seated at a distance of around 60 cm in a laboratory environment with normal indoor lighting. The subjects are provided with detailed evaluation criteria and trained before the formal experiment to be familiar with the experiment. Individuals who demonstrate low agreement accuracy are excluded from the experiment. Finally, we collect 84 (28 subjects × 3 perspectives) ratings and an average of 7 preference descriptions for each AIGI, resulting in a total number of 201,600 (28 × 3 × 2400) quality ratings and 16,800 (7 × 2,400) detailed explanations.
@@ III-D Question Answering Pair Generation
In order to enable the detailed human preference explanation ability for an IQA model, we collect numerous question-answering pairs tailored for instruction tuning. Some of the questions are consistent with the setup questions in our subjective experiment, and some are extended based on the degradation or quality-related labels. The extended question-answering pairs are generated with the help of large language models (LLMs). Specifically, we feed the collected quality labels and detailed explanations into LLMs, then prompt LLMs to ask questions for these answers. This step can greatly expand the scale of VQA pairs. Fig. 5 fine-grained human preference annotations, we can improve the perceptual explanation capability of LLMs and further improve the performance of IQA models.
@@ III-E Human Annotation Analysis
We first analyze the subjective quality rating and MOS distribution from the perspectives of quality, authenticity, text-image correspondence respectively. As shown in Fig. 6, the images in AIGCIQA2023+ database cover a wide range of perceptual quality. We further analyze the distributions of fine-grained annotations. Fig. 7 demonstrates the distribution of fine-grained annotations for all models, and Fig. 8 further shows the distribution of fine-grained annotations for different generation models including Glide [44], Lafite [45], Unidiffuser [3], Stable-diffusion [47], Controlnet [48], DALLE [46]. It can be observed that the performance of DALLE [46] is the best among all the six AIGI models. The performance of Glide [44] and Lafite [45] is relatively poor compared to other models in terms of the fine-grained assessment. Besides, Lafite [45] is better than Glide [44] in terms of text-image consistency, but far worse than Glide [44] in terms of the authenticity perspective. This also manifests that different AIGI models have different defects, which further illustrates the importance of evaluating AIGIs from multiple perspectives. And our analysis can help users choose appropriate AIGI models according to the different preference emphasis.
@@ IV-A1 Prompt Segmentation
The content and length of the input prompt are of great significance in the text-image correspondence performance evaluation. The information provided by the original prompt may not be cohesive or meaningful enough, thus may lead to poor performance on the evaluation of AIGIs, particularly in terms of text-image correspondence. To address this issue, we apply a “raw prompt + LLM-refined prompt” mechanism to make the text more understandable for text models. Specifically, we utilize a LLM to transform text prompts into three task-specific annotations including “style”, “content”, and “atmosphere”. “Style” words describe the whole image style, including painting, artistic, realistic, fashion, texture, fiction, etc. If the original input text does not have style cues, the default style is “realistic”. “Content” words are nouns with adjectives describing the image content. “Atmosphere” words refer to the emotional and psychological elements associated with the image, including mood and feeling conveyed by the scene. All of these three items lying after the raw prompt are crucial to the assessment of text-image correspondence.
@@ IV-C Training and Fine-tuning Strategy
Method SRCC PLCC KRCC SRCC PLCC KRCC SRCC PLCC KRCC SRCC PLCC KRCC
@@ IV-C1 Score Regression Pre-training
$&\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left|Q_{\text{predict}}(i)-Q_{\text{label}}(i)\right|\end{aligned}\vspace{-3mm}$ (4)
@@
$\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(T,P_{b},P_{w})\sim\mathcal{D}}[\mathop{\log}(\sigma(f_{\theta}(T,P_{b})-f_{\theta}(T,P_{w})))]\end{split}$ (5)
@@ IV-C2 Vision-to-Language Instruction Tuning
IV-C3 Feedback from the explanation module to evaluating module
To further strengthen the connection between the two frameworks and help predict more accurate scores, we take the instruction-aware representations from the second framework as a reward. As shown in Fig. 10, we concanate the instruction-aware representations and the quality-aware representations as the input for multi-head quality regression, and then fine-tune the three regressors respectively with corresponding instructions. For example, when finetuning the quality regressor, the instruction is “how is the quality of the image”. The instruction-aware representations according to this instruction are highly quality-aware. Consequently, taking these representations from the second framework as a reward to the first framework leads to more accurate quality score prediction. As shown in Table XIV, the model gains better performance through the feedback.
V Experiment
V-A Experimental Protocol
TABLE III: Perception metric performance comparisons on the AGIQA-3K database. The best results are marked in RED and the second results are marked in BLUE.
Method SRCC PLCC KRCC
CEIQ [63] 0.3228 0.4166 0.2220
@@ V-A Experimental Protocol
MINT-IQA (Ours) 0.8919 0.9328 0.7224
TABLE IV: Alignment metric performance comparisons on the AGIQA-3K database. The best results are marked in RED and the second results are marked in BLUE.
Method SRCC PLCC KRCC
CLIP [20] 0.5972 0.6839 0.4591
@@
MINT-IQA (Ours) 0.8749 0.9282 0.7036
Some baseline results are evaluated by ourselves. For traditional handcrafted-based models, they are directly evaluated on the corresponding databases. For vision-language pre-training models, we simply load the pre-trained weights for inference. CLIP score and BLIP score are calculated directly as cosine similarity between text and image embeddings. For deep learning-based models, we use the same training and testing split as the previous literature. For our proposed model, the number of regressors in our model is flexible, corresponding to the varying number of evaluation dimensions in the test databases. Since traditional IQA datasbases have only one dimension for assessment, only one regressor is used for score prediction. The models are implemented with PyTorch and trained on a 40GB NVIDIA RTX A6000 GPU with batch size of 8 on AIGCIQA2023+ database and batch size 16 on the other five databases. The initial learning rate is set to 1e-5, and decreased using the cosine annealing strategy. We employ Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
V-B Performance on AIGC IQA Databases
We evaluate the performance of our proposed metric, MINT-IQA, on three recently released AIGC IQA databases including AIGCIQA2023 [17], AGIQA-3K [16], and ImageReward [15]. The performance comparison results are demonstrated in Tables II-V. Handcrafted-based models show poor performance on AIGC IQA databases, indicating that the handcrafted features mainly targeted for natural images are ineffective for evaluating AIGIs. On the other hand, deep learning-based methods achieve relatively better results. However, they are still far away from satisfactory.
MINT-IQA attains state-of-the-art performance across all three AIGC IQA datasetbases, substantially outperforming current IQA models. As shown in Table II, our proposed MINT-IQA model achieves the best performance on our AIGCIQA2023+ from all three perspectives, which demonstrates the effectiveness of MINT-IQA in evaluating human preferences from multiple perspectives. Additionally, as demonstrated in Table III and Table IV, our model outperforms other methods in measuring the quality perception and text-image alignment on AGIQA-3K [16]. This highlights the versatility of our model in understanding human preference for AIGIs in terms of both quality perception perspective and alignment perspective.
Our proposed method also shows superior performance on the ImageReward dataset [15] compared to the ImageReward model, as shown in Table V, even though we change the training data format from a single image and its MOS to a pair of images and their comparisons. This further manifests the generalization ability of our model.
TABLE V: Human preference prediction metric performance comparisons on the ImageReward database. Preference accuracy is conducted based on the test set with 466 prompts (6,399 comparisons). The best results are marked in RED and the second results are marked in BLUE.

### v2

# Quality Assessment for AI Generated Images with Instruction Tuning

@@ Abstract
Artificial Intelligence Generated Content (AIGC) has grown rapidly in recent years, among which AI-based image generation has gained widespread attention due to its efficient and imaginative image creation ability. However, AI-generated Images (AIGIs) may not satisfy human preferences due to their unique distortions, which highlights the necessity to understand and evaluate human preferences for AIGIs. To this end, in this paper, we first establish a novel Image Quality Assessment (IQA) database for AIGIs, termed AIGCIQA2023+, which provides human visual preference scores and detailed preference explanations from three perspectives including quality, authenticity, and correspondence. Then, based on the constructed AIGCIQA2023+ database, this paper presents a MINT-IQA model to evaluate and explain human preferences for AIGIs from Multi-perspectives with INstruction Tuning. Specifically, the MINT-IQA model first learn and evaluate human preferences for AI-generated Images from multi-perspectives, then via the vision-language instruction tuning strategy, MINT-IQA attains powerful understanding and explanation ability for human visual preference on AIGIs, which can be used for feedback to further improve the assessment capabilities. Extensive experimental results demonstrate that the proposed MINT-IQA model achieves state-of-the-art performance in understanding and evaluating human visual preferences for AIGIs, and the proposed model also achieves competing results on traditional IQA tasks compared with state-of-the-art IQA models. The AIGCIQA2023+ database and MINT-IQA model are available at: https://github.com/IntMeGroup/MINT-IQA
@@ I Introduction
Artificial Intelligence Generated Content (AIGC) refers to the content generated with the assistance of AI, including texts, images, audios, etc. As an important part of AIGC, AI Generated Images (AIGIs) have gained significant attention due to their broad application prospects, thus, many AI-based image generation methods have been proposed in recent years, such as DALLE [1], Stable-diffusion [2], Unidiffuser [3], etc. However, due to the limitations of model capacity, computing overhead, and human-AI interaction gap, etc., not all generated images can satisfy users’ requirements, and many AIGIs do not conform to human visual preferences or expectations. Unlike common image content, such as Natural Scene Images (NSIs) [4, 5], Screen Content Images (SCIs) [6, 7] etc., which generally encounters common degradations such as noise, blur, compression [8, 9, 10, 11, 12, 13, 14, 15, 16, 17], AIGIs exhibit unique distortions including unreal structures, unreasonable component combinations, etc. [18, 19, 20] Moreover, for text-to-image generation methods, the generated images may not accurately correspond to the semantics of the text prompts or may not meet the expectations of users [21].
@@
To address the aforementioned challenges, in this paper, we first extend the previous AIGCIQA2023 database [25] to AIGCIQA2023+, towards better understanding and evaluating human preferences for AIGIs, and then propose a novel method, termed MINT-IQA, towards automatically evaluating and explaining the visual preference and expectation attributes for AIGIs from Multi-perspectives with INstruction Tuning. As an extension version of the conference version [25], the differences rely on two aspects: (1) annotation extension and (2) a new method design. Specifically, as shown in Fig. 1, the established AIGCIQA2023+ contains 2400 AIGIs generated using 6 AIGI models based on 100 text prompts, 7,200 Mean Opinion Scores (MOSs) and 16,800 detailed preference-explanation annotations for these AIGIs. The evaluation perspectives include quality, authenticity, and text-image correspondence. We further propose a MINT-IQA model, which allows for a more comprehensive evaluation of the human preferences for AIGIs from multi-perspectives, resulting in state-of-the-art performance on AIGC IQA databases. To further enhance the degradation understanding and explanation capabilities of our model, we adopt the vision-language instruction tuning strategy, and fine-tune the MINT-IQA model based on the AIGCIQA2023+ database. This approach effectively leverages the frozen Large Language Model (LLM) to provide detailed explanations of the human preference basis for each AIGI as shown in Fig. 2, which can be used for feedback to further improve the assessment capabilities as shown in Section V. Additionally, MINT-IQA model also achieves the best performance on traditional IQA databases, which further demonstrates its generality in understanding and evaluating human preferences.
@@
  • We establish an AI-generated image quality assessment database, termed AIGCIQA2023+, to conduct comprehensive understanding and evaluation research on human visual preferences for AIGIs.
@@ II-A Image Quality Assessment Databases
Traditional IQA Databases LIVE [30] MOS Overall 779 25,000 No
CSIQ [31] MOS Overall 866 5,000 No
TID2013 [32] MOS Overall 3,000 524,340 No
CLIVE [33] MOS Overall 1,162 350,000 No
KADID [34] MOS Overall 10,125 303,750 No
@@
Many IQA databases have been established in the literature [39, 40, 41, 42, 43]. As shown in Table I, according to the image acquisition method, they can be categorized into two types: (1) traditional IQA databases and (2) AIGC IQA databases. Traditional IQA databases consist of natural images with realistic camera distortions such as noise, blur, compression, etc. LIVE [30], CSIQ [31], TID2013 [32], and CLIVE [33] are fundamental traditional IQA databases and have been widely used for developing IQA models. KADID [34], KonIQ-10k [44] and SPAQ [36] expand the data scale and cover a wide range of scene categories. AVA [37] is a large-scale database constructed for aesthetic image quality assessment. Recently, the development of text-to-image generative models has led to an explosion of AI-generated images. Thus many AIGC IQA databases have also been constructed, which contain generated images with unique distortions including unreal structures, unreasonable component combinations, etc. DiffusionDB [38] is a large-scale database containing over 1.8 million text-image pairs without score-specific annotations. Pick-A-Pic [22] and HPS [18] provide subjective annotations of human preferences by selecting the most preferred one among a group or a pair of AIGIs. ImageReward [23] contains preference ranking annotations for AIGIs. The aforementioned AIGC IQA databases only evaluate the AIGIs from the overall dimension, which cannot reflect and characterize complex human visual preferences for AIGIs. AGIQA-3K [24] refines the rating granularity by scoring two dimensions including perception and alignment. Our preliminary AIGCIQA2023 database [25] contains subjective quality ratings from three dimensions including quality, authenticity, and correspondence. However, the aforementioned works lack detailed annotations of the human preference for AIGIs, which motivates us to further extend the AIGCIQA2023 database to include language annotations.
@@ II-B Image Quality Assessment Models
Numerous image and video quality assessment models have been proposed in the literature [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], which aims to objectively and quantitatively evaluate the perceptual quality. Based on different feature extraction and evaluation methods, current popular models for image quality assessment can be classified into three categories: (1) Handcrafted-based models, including: NIQE [60], BRISQUE [61], QAC [62], BMPRI [63], HOSA [64], BPRI [65], HIGRADE [66], etc. These models generally extract features based on prior knowledge related to image quality. (2) Deep learning-based models, including: CNNIQA [67], DBCNN [68], WaDIQaM-NR [69], StairIQA [70], CLIPAGIQA [45], Q-Bench [48], etc. These models characterize quality-aware information by training deep neural networks using labeled data. (3) With the popularity of text-to-image generation, many vision-language pre-training models have been adopted to evaluate the text-image alignment, including: CLIP [28], BLIP [29], FLIP [71], Aesthetic [72], HPS [18], PickScore [22], ImageReward [23], StairReward [24], etc. These models are often used for evaluating generated images based on its text prompt considering a mixture of elements such as text-image alignment, fidelity, aesthetics, etc. However, the aforementioned IQA models can only evaluate a single-dimension preference for AI-generated images or videos, which lacks the ability to conduct a comprehensive assessment from multiple perspectives. This limitation motivates us to propose a unified model to evaluate AIGIs from multiple dimensions.
@@ III-A AIGI Collection
We adopt six latest text-to-image generative models, including Glide[73], Lafite[74], DALLE[75], Stable-diffusion[76], Unidiffuser[3], Controlnet[77], to produce AIGIs by using open source code and default weights. To ensure content diversity and catch up with the practical application requirements, we collect diverse texts from the PartiPrompts website [78] as the prompts for AI-based image generation. The text prompts can be simple, allowing generative models to produce imaginative results. They can also be complex, which raises the challenge for generative models. We select 10 scene categories from the prompt set, including people, animals, artifacts, illustrations, indoor scenes, outdoor scenes, vehicles, produce & plants, food & beverage, and world knowledge. Each scene contains 10 challenge categories, including basic, simple detail, fine-grained detail, complex, quantity, imagination, style & format, perspective, writing & symbols, and linguistic structures. Overall, we collect 100 text prompts (10 scene categories $\times$ 10 challenge categories) from PartiPrompts[78]. Based on the selected prompts, we generate 4 different images for each generative model. Therefore, the constructed database totally contains 2400 AIGIs (4 images $\times$ 6 models $\times$ 100 prompts).
@@ III-B Subjective Experiment Setup
The AIGIs are presented in a random order on an iMac monitor with a resolution of 4096 × 2304, using an interface designed with Python Tkinter as shown in Fig. 4. The interface allows viewers to browse the previous and next AIGIs, and evaluate them by scoring within the range from 0 to 5 as well as providing fine-grained explanations. A total of 76 graduate students participate in the experiment. The subjects are seated at a distance of around 60 cm in a laboratory environment with normal indoor lighting. To ensure consistency, we have developed and documented the annotation process. For the purpose of explaining human preferences, we have also established criteria for providing explanations in a standardized format, and have included numerous illustrative examples. The subjects are provided with detailed evaluation criteria and trained before the formal experiment to be familiar with the experiment. Individuals who demonstrate low agreement accuracy are excluded from the experiment. Finally, we collect 84 (28 subjects × 3 perspectives) ratings and an average of 7 preference descriptions for each AIGI, resulting in a total number of 201,600 (28 × 3 × 2400) quality ratings and 16,800 (7 × 2,400) detailed explanations.
@@ III-D Question Answering Pair Generation
In order to enable the detailed human preference explanation ability for an IQA model, we collect numerous question-answering pairs tailored for instruction tuning. Some of the questions are consistent with the setup questions in our subjective experiment, and some are extended based on the degradation or quality-related labels. The extended question-answering pairs are generated with the help of LLM (LLaMA[79]). Specifically, we feed the collected quality labels and detailed explanations into LLM, then prompt LLM to ask questions for these answers. This step can greatly expand the scale of VQA pairs. Fig. 5 fine-grained human preference annotations, we can improve the perceptual explanation capability of LLMs and further improve the performance of IQA models.
@@ III-E Human Annotation Analysis
We first analyze the subjective quality rating and MOS distribution from the perspectives of quality, authenticity, text-image correspondence respectively. As shown in Fig. 6, the images in AIGCIQA2023+ database cover a wide range of perceptual quality. We further analyze the distributions of fine-grained annotations. Fig. 7 demonstrates the distribution of fine-grained annotations for all models, and Fig. 8 further shows the distribution of fine-grained annotations for different generation models. It can be observed that the performance of DALLE [75] is the best among all the six AIGI models. The performance of Glide [73] and Lafite [74] is relatively poor compared to other models in terms of the fine-grained assessment. Besides, Lafite [74] is better than Glide [73] in terms of text-image consistency, but far worse than Glide [73] in terms of the authenticity perspective. This also manifests that different AIGI models have different defects, which further illustrates the importance of evaluating AIGIs from multiple perspectives. And our analysis can help users choose appropriate AIGI models according to the different preference emphasis.
@@ IV-A1 Prompt Segmentation
The content and length of the input prompt are of great significance in the text-image correspondence performance evaluation. The information provided by the original prompt may not be cohesive or meaningful enough, thus may lead to poor performance on the evaluation of AIGIs, particularly in terms of text-image correspondence. To address this issue, we apply a “raw prompt + LLM-refined prompt” mechanism to make the text more understandable for text models. Specifically, we utilize a LLM (LLaMA [79]) to transform text prompts into three task-specific annotations including “style”, “content”, and “atmosphere”. “Style” words describe the whole image style, including painting, artistic, realistic, fashion, texture, fiction, etc. If the original input text does not have style cues, the default style is “realistic”. “Content” words are nouns with adjectives describing the image content. “Atmosphere” words refer to the emotional and psychological elements associated with the image, including mood and feeling conveyed by the scene. All of these three items lying after the raw prompt are crucial to the assessment of text-image correspondence.
@@ IV-C Training and Fine-tuning Strategy
SRCC PLCC KRCC SRCC PLCC KRCC SRCC PLCC KRCC SRCC PLCC KRCC
@@ IV-C1 Score Regression Pre-training
$&\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left|Q_{\text{predict}}(i)-Q_{\text{label}}(i)\right|,\end{aligned}\vspace{-3mm}$ (4)
@@
$\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(T,P_{b},P_{w})\sim\mathcal{D}}[\mathop{\log}(\sigma(f_{\theta}(T,P_{b})-f_{\theta}(T,P_{w})))],\end{split}$ (5)
@@ IV-C2 Vision-to-Language Instruction Tuning
TABLE III: Performance comparisons on the AGIQA-3K database. The best results are marked in RED and the second results are marked in blue.
Dimension Method SRCC PLCC KRCC
Quality CEIQ [91] 0.3228 0.4166 0.2220
@@
CLIP-IQA+ [90] 0.8428 0.8879 0.6556
CLIP-AGIQA [45] 0.8747 0.9190 0.6976
[…]
Improvement + 1.7% + 1.4% + 2.5%
Alignment CLIP [28] 0.5972 0.6839 0.4591
@@
[…]
Improvement + 12.8% + 7.5% + 14.8%
IV-C3 Feedback from the explanation module to evaluating module
To further strengthen the connection between the two frameworks and help predict more accurate scores, we take the instruction-aware representations from the second framework as a reward. As shown in Fig. 10, we concanate the instruction-aware representations and the quality-aware representations as the input for multi-head quality regression, and then fine-tune the three regressors respectively with corresponding instructions. For example, when finetuning the quality regressor, the instruction is “how is the quality of the image”. The instruction-aware representations according to this instruction are highly quality-aware. Consequently, taking these representations from the second framework as a reward to the first framework leads to more accurate quality score prediction. As shown in Table XIII, the model gains better performance through the feedback.
V Experiment
V-A Experimental Protocol
TABLE IV: Human preference prediction metric performance comparisons on the ImageReward database. Preference accuracy is conducted based on the test set with 466 prompts (6,399 comparisons). The best results are marked in RED and the second results are marked in blue.